# Sentiment-Aware Stock Price Prediction with Transformer and LLM-Generated Formulaic Alpha

### v1

@@ 1 Introduction
Deep learning models, especially Transformers, have recently shown strong results in sequence modeling tasks. Their ability to learn long-range dependencies makes them suitable for predicting price movements [1, 2, 3]. However, using only raw financial features often leads to noisy and unstable predictions. To improve model performance, many researchers and practitioners create special features called "alpha factors" or "alphas." These are formulas designed to capture patterns that can predict future returns. Equation 1 is an example of a formulaic alpha.
@@
In this paper, we propose a new framework that uses prompt-based LLMs to automatically generate formulaic alphas from stock features, technical indicators, and sentiment scores. These generated alphas serve as additional inputs to a vanilla Transformer model that predicts next-day stock prices. By introducing new machine-generated alphas, it is expected that alpha decay will slow down and stronger predictive signals will be maintained over time.
@@ 2.1 Stock Price Prediction with Machine Learning Models
Transformers, introduced later, overcame these limitations by relying on self-attention mechanisms. Transformers can capture long-term dependencies more effectively and have become a popular choice for time series forecasting tasks, including in finance. Several studies have adapted the Transformer architecture for stock prediction by treating historical prices, technical indicators, and other features as sequential inputs [16, 17]. Some approaches modify the attention mechanism to better focus on important market events [18]. Others combine Transformers with other models, such as Convolutional Neural Networks (CNNs), to capture both short-term and long-term patterns [19, 20]. The flexibility of Transformers in handling different types of input data makes them attractive for financial forecasting.
@@ 3.1 Data
The stock data for five companies, Apple, HSBC, Pepsi, Toyota, and Tencent, is collected using the yfinance Python library. The data cover the period from May 1, 2016, to May 8, 2024. These companies are selected to include a mix of industries and markets. The stock data includes Open, High, Low, Close, and Volume information.
@@
In addition to stock data, daily news articles related to the five companies as well as their related companies are downloaded from websites such as Yahoo News using the Financial News Feed and Stock News Sentiment Data API [^1] https://eodhd.com/financial-apis/stock-market-financial-news-api. Based on our previous study, related companies are defined as those that appear frequently in news articles about the five target companies [6]. For example, BYD appears in Toyota’s news articles over one hundred times, suggesting a connection between the two companies. The news articles help capture market sentiment, which can have an important impact on stock prices. Previously, we used NER to extract entity names from news articles related to the target companies, based on their frequency of appearance. All identified listed companies were then ranked from highest to lowest frequency, and for each target company, the top ten most frequently co-mentioned companies were selected as related companies. Table 2 shows the target companies and their corresponding related companies.
@@
HSBC JPMorgan, Standard Chartered, Morgan Stanley, UBS, HSBC, Citigroup, Goldman Sachs, Credit Suisse, NatWest, Apple, Microsoft
Infusystem Northwest Pipe, Citizens & Northern, Cabaletta Bio, Luokung Technology, Sanara MedTech, Adicet Bio, Westamerica, CECO Environmental
@@
2. Technical indicators (e.g., RSI, moving averages, MACE),
@@
The first 70% of the data is used for training and the remaining 30% of the data is used for testing. A sliding window of 5 days is applied to the input features.
@@ 3.2 Formulaic Alpha Generation Using LLM
To generate new trading formulas, an LLM called deepseek-r1-distill-llama-70b is used. This model is developed by the American company Groq [^2] https://groq.com/. It is a distilled version of a larger model. Knowledge distillation is a technique in which a smaller model (student) learns to mimic the behavior of a larger, more powerful model (teacher). This process reduces the size and computational cost of the model while keeping much of the original performance.
@@ 3.3 Stock Prediction Using Transformer
The features, the LLM-generated alphas, are fed into an Encoder-Decoder Transformer model (Figure 2). The Transformer is designed to predict the closing price the next day for each stock.
@@
The model is trained using the Mean Squared Error (MSE) loss function between the predicted and true closing prices. The model is optimized using the Adam optimizer with a learning rate of $5\times 10^{-5}$ over 50 epochs. The batch size varies by stock, ranging from 8 to 64. Key hyperparameters, such as the learning rate and batch size, are selected through empirical tuning based on the validation performance. A grid of commonly used values is explored, and the final configuration is chosen to balance predictive accuracy and training stability.
@@ 3.4 Models Used for Performance Comparison
In this study, some commonly used models in stock forecasting are added for comparison. The following is a brief description of each model used in this study.
  • Long Short-Term Memory (LSTM) is a type of recurrent neural network that is good at learning patterns in time series data. It uses memory cells and gates to keep important information and forget less useful signals over time.
  • Temporal Convolutional Network (TCN) is a deep learning model that handles sequence data using 1D convolution layers. Unlike RNNs, it processes sequences in parallel and captures long-range dependencies using dilated convolutions.
  • Support Vector Regression (SVR) is a machine learning method that finds a line or curve to fit the data while keeping errors within a set margin. It works well on small to medium-sized datasets and handles both linear and non-linear trends.
  • Random Forest is an ensemble model that builds many decision trees and combines their results. It reduces overfitting and improves prediction accuracy by averaging the outcomes of individual trees.
@@ 3.5 Evaluation
Furthermore, to highlight the novelty of incorporating sentiments from both target companies and their related firms, we evaluate the impact of integrating sentiment information into a Transformer model for stock price forecasting, which will be illustrated in Section 4.3. For both the baseline scenario without sentiment and the experimental scenario with sentiment, we conduct ten independent runs per stock. This number is chosen as a common practice in empirical machine learning research to balance statistical robustness with computational efficiency. We report the average MSE and the standard deviation in each case to capture performance and account for training variability. This ensures a fair and statistically sound comparison, allowing us to assess not only the average performance improvement of sentiment features, but also the consistency of the model predictions across multiple runs.
@@ 4 Results
Table 4 shows the LLM-generated formulaic alphas for all the five companies. Table 6 demonstrates the Transformer performance with and without sentiment scores as input. Figure 3 shows the reasoning process of the LLM to generate the formulaic alphas for Toyota. Table 5 shows the prediction performance in terms of MSE for all companies using the formulaic alphas in Table 4. Figure 4 visualizes the prediction performances of all models for HSBC stock prediction. Based on the figures, it is obvious that all models can capture the stock movement accurately when using the LLM-generated features as input. More details of the results will be illustrated in the subsequent subsections.
@@ 4.2 Stock Prediction Performance
According to Table 5, for most of the time, Transformer has the best performance in terms of MSE. The plotted graph for the Transformer (Figure 4) shows that the predicted stock prices closely follow the actual stock price trend, indicating its strong ability to capture underlying patterns in the data.
@@
For traditional machine learning models, the SVR has the worst performance. The random forest ensemble method can sometimes generate competitive performance compared to the Transformer (e.g., HSBC and Toyota).
@@
HSBC alpha1_t = (C_t - O_t) / O_t + 0.3 * Apple_polarity_t + 0.2 * (SMA_5_t - SMA_20_t) alpha2_t = (RSI_14_t / 14) * Momentum_10_t + 0.4 * UBS_polarity_t - 0.2 * BB_Lower_t alpha3_t = (MACD_t - MACD_Signal_t) + 0.5 * Microsoft_polarity_t + 0.2 * (V_t / V_t_1) alpha4_t = (C_t / SMA_20_t) - 1 + 0.3 * Morgan_Stanley_polarity_t - 0.1 * (BB_Upper_t - BB_Lower_t) alpha5_t = (Momentum_3_t + Momentum_10_t) / 2 + 0.4 * Goldman_Sachs_polarity_t - 0.2 * OBV_t
Pepsi alpha1_t = Momentum_3_t + Amazon_polarity_t alpha2_t = (SMA_5_t - SMA_20_t) + Microsoft_polarity_t alpha3_t = (V_t / V_t_minus_1) * Walmart_polarity_t alpha4_t = (MACD_t - MACD_Signal_t) + IBM_polarity_t alpha5_t = (50 - RSI_14_t) + TSLA_polarity_t
Tencent alpha1_t = (Momentum_3_t + Momentum_10_t) / 2 + 0.3 * (V_t / V_t-1) + 0.2 * polarity_avg alpha2_t = (SMA_5_t - SMA_20_t) + 0.5 * (C_t - O_t) / O_t + 0.4 * polarity_avg alpha3_t = (RSI_14_t - 50) / 50 + 0.6 * (OBV_t - OBV_t-20) / OBV_t-20 + 0.3 * polarity_avg alpha4_t = (Momentum_3_t + Momentum_10_t) / 2 + 0.4 * (EMA_10_t - EMA_10_t-1) + 0.2 * polarity_avg alpha5_t = (MACD_t - MACD_Signal_t) + 0.5 * (BB_Upper_t - BB_Lower_t) / BB_Lower_t + 0.3 * polarity_avg # polarity_avg = mean of 12 tech company polarities
Toyota alpha1_t = (Momentum_3_t + Momentum_10_t) / 2 + (Toyota_polarity_t + Nissan_polarity_t + Honda_polarity_t) / 3 alpha2_t = RSI_14_t * (1 if RSI_14_t < 30 else 0) * (1 if Toyota_polarity_t > 0 else 0) alpha3_t = 1 if (SMA_5_t > SMA_20_t) and (Toyota_polarity_t > 0) else 0 alpha4_t = (C_t - BB_Lower_t) / (BB_Upper_t - BB_Lower_t) * (1 if Apple_polarity_t > 0 and Amazon_polarity_t > 0 else 0) alpha5_t = (OBV_t > OBV_t-1) * (C_t > C_t-1) * (1 if BYD_polarity_t > 0 and TSLA_polarity_t > 0 else 0)
@@
Table 5: MSE Comparison of Models for Stock Price Prediction with LLM-Generated Alphas

### v2

@@ 1 Introduction
Deep learning models, especially Transformers, have recently shown strong results in sequence modeling tasks. Their ability to learn long-range dependencies makes them suitable for predicting price movements [20, 49, 46, 12]. However, using only raw financial features often leads to noisy and unstable predictions. To improve model performance, many researchers and practitioners create special features called "alpha factors" or "alphas." These are formulas designed to capture patterns that can predict future returns [7]. Equation 1 is an example of a formulaic alpha.
@@
In this paper, we propose a new framework that uses prompt-based LLMs to automatically generate formulaic alphas from stock features, technical indicators, and sentiment scores. These generated alphas serve as additional inputs to a vanilla Transformer model (encoder-decoder structure) [44] that predicts next-day stock prices. By introducing new machine-generated alphas, it is expected that alpha decay will slow down and stronger predictive signals will be maintained over time.
@@ 2.1 Stock Price Prediction with Machine Learning Models
Transformers, introduced later, overcame these limitations by relying on self-attention mechanisms. Transformers can capture long-term dependencies more effectively and have become a popular choice for time series forecasting tasks, including in finance. Several studies have adapted the Transformer architecture for stock prediction by treating historical prices, technical indicators, and other features as sequential inputs [6, 27, 10, 9]. Some approaches modify the attention mechanism to better focus on important market events [23]. Others combine Transformers with other models, such as Convolutional Neural Networks (CNNs), to capture both short-term and long-term patterns [52, 28]. The flexibility of Transformers in handling different types of input data makes them attractive for financial forecasting.
@@ 3.1 Data
Following common practice in related stock prediction studies, which often evaluate models using a limited set of representative stocks (typically between one to ten) [43, 45, 1, 4, 5, 13], this study selects five companies for empirical analysis. This choice allows for a focused evaluation of the proposed method while maintaining manageable computational complexity. The stock data for five companies, Apple, HSBC, Pepsi, Toyota, and Tencent, is collected using the yfinance Python library. The data cover the period from May 1, 2016, to May 8, 2024. These companies are chosen because they operate in different industries and markets, including technology, finance, consumer goods, automotive, and internet services, which allows the proposed framework to be evaluated under diverse market characteristics. All five stocks are highly liquid and widely followed, ensuring data availability. The stock data includes Open, High, Low, Close, and Volume information.
@@
In addition to stock data, daily news articles related to the five companies as well as their related companies are downloaded from websites such as Yahoo News using the Financial News Feed and Stock News Sentiment Data API [^1] https://eodhd.com/financial-apis/stock-market-financial-news-api. All news is in English. Based on our previous study, related companies are defined as those that appear frequently in news articles about the five target companies [8]. For example, BYD appears in Toyota’s news articles over one hundred times, suggesting a connection between the two companies. The news articles help capture market sentiment, which can have an important impact on stock prices. Previously, we used NER to extract entity names from news articles related to the target companies, based on their frequency of appearance. All identified listed companies were then ranked from highest to lowest frequency, and for each target company, the top ten most frequently co-mentioned companies were selected as related companies. The Python library spaCy is used to perform NER. Figure 2 demonstrates the whole process of the related company selection. Table 2 shows the target companies and their corresponding related companies. Furthermore, the related companies were extracted using the whole news dataset. However, the extraction step is performed once and produces a fixed set of related companies that is independent of the model training and prediction process. During training and testing, the model only uses news and price information from the corresponding time periods, and no test-period news content is used to update model parameters or input features. As a result, future information is not exploited in the predictive task. Figure 3 shows the related company extraction and model training procedure.
News ArticlesNamed EntityRecognition (NER)Extracted CompanyNamesCo-occurrenceFrequency CountRelated CompanySelectionIdentify organization entitiesCount joint appearances within the news article
Figure 2: Company extraction using named entity recognition and related company selection based on co-occurrence frequency. [8]
Daily News DatasetRelated-Company Extraction (NER)Fixed Related-Company Lists Example List: Apple Amazon Meta GoogleTemporal SplitTraining / TestingModel Training(with training period only) Example: Apple_polarity: 0.15 Amazon_polarity: 0.56 Meta_polarity: -0.8 Google_polarity: 0 (Google is absent from the news)Model Evaluation(with testing period only) Example: Apple_polarity: 0.25 Amazon_polarity: 0.06 Meta_polarity: -0.28 Google_polarity: 0.17 (Google appears in the news) Related-company lists are extracted once and fixed
Figure 3: Illustration of related-company extraction and model training procedure without data leakage. The related-company list is extracted once and kept fixed. During model training, if a company does not appear in the training-period news (e.g., Google), its polarity value is set to zero, which will not cause information leakage during training.
@@
Algorithm 1 Time-Series Stock Prediction Pipeline
Input: Daily news and stock dataset $D$, sliding window size $w$, prediction horizon $h$
Output: Trained model and test predictions
 Split $D$ chronologically into training set $D_{train}$ and testing set $D_{test}$
 Extract related-company lists from $D$ (fixed, no update during training/testing)
 Compute LLM-generated features for all companies
 Fit MinMax scalers on $D_{train}$ features and labels
 Apply scalers to $D_{train}$ and $D_{test}$ (transform only)
 Construct sliding windows of length $w$:
for each time $t$ in $D_{train}$ do
  $X_{train}[t]\leftarrow$ features from twt-w to t1t-1
  $y_{train}[t]\leftarrow$ close price at $t+h$
end for
for each time $t$ in $D_{test}$ do
  $X_{test}[t]\leftarrow$ features from twt-w to t1t-1
  $y_{test}[t]\leftarrow$ close price at $t+h$
end for
 Train prediction model (e.g., Transformer, TCN) on $(X_{train},y_{train})$
 Evaluate model on $(X_{test},y_{test})$
return Trained model and test predictions
@@
HSBC JPMorgan, Standard Chartered, Morgan Stanley, UBS, Citigroup, Goldman Sachs, Credit Suisse, NatWest, Apple, Microsoft
@@
2. Technical indicators (e.g., RSI, moving averages, MACD),
@@
All timestamps for the stock and news data are recorded in UTC to avoid the time zone inconsistency between different stocks. The news data used in this study are at the daily level, and intraday timestamps are not used. Each day’s polarity score is computed solely from news published on that calendar day.
Following the temporal order of the time series, the first 70% of the data is used for training and the remaining 30% of the data is used for testing. A sliding window of 5 days is applied to the input features. The stock prediction pipeline is illustrated in Algorithm 1.
@@ 3.2 Formulaic Alpha Generation Using LLM
To generate new alpha formulas, we use the deepseek-r1-distill-llama-70b model, which is a distilled variant of the DeepSeek-R1 model based on the LLaMA architecture and is provided by Groq [^2] https://groq.com/ for inference and deployment. Knowledge distillation is a technique in which a smaller model (student) learns to mimic the behavior of a larger, more powerful model (teacher). This process reduces the size and computational cost of the model while keeping much of the original performance.
@@
All LLM-generated formulaic alphas are normalized using Min–Max scaling before being fed into the prediction models. The scaling parameters are estimated from the training set and then applied to the testing set to avoid information leakage.
@@ 3.3 Stock Prediction Using Transformer
The features, the LLM-generated alphas, are fed into a vanilla Transformer model (Figure 4). The Transformer is designed to predict the closing price the next day for each stock.
@@
The model is trained using the Mean Squared Error (MSE) loss function between the predicted and true closing prices. The model is optimized using the Adam optimizer with a learning rate of $5\times 10^{-5}$ over 50 epochs. The batch size varies by stock, ranging from 3 to 64. Key hyperparameters, such as the learning rate and batch size, are selected through empirical tuning based on the training performance. All hyperparameter tuning was conducted exclusively on the training set. The test set was held out and used only for final evaluation. The adoption of a small batch size and a low learning rate serves to reduce the risk of overfitting. As demonstrated in prior studies [26, 29, 36], these training constraints help the model capture generalizable patterns rather than idiosyncratic noise in the training data. A grid of commonly used values is explored, and the final configuration is chosen to balance predictive accuracy and training stability. Although the Transformer model was tuned using a simpler method, it still achieved competitive performance, indicating robustness against overfitting. Table 4 shows the final hyperparameters settings of the Transformer model.
@@
Table 4: Final hyperparameter settings of the Transformer model
Hyperparameter Value
Encoder layers 3
Decoder layers 2
Hidden size 512
Feed-forward size 512
Embedding size 512
Dropout rate 0.0
Activation function GELU
Window length 5
@@ 3.4 Models Used for Performance Comparison
In this study, some commonly used models in stock forecasting are added for comparison. The following is a brief description of each model used in this study. We employ Optuna [2] to automatically search for suitable hyperparameter configurations for deep learning models and GridSearchCV for machine learning models such as XGBoost.
  • Long Short-Term Memory (LSTM) is a type of recurrent neural network that is good at learning patterns in time series data. It uses memory cells and gates to keep important information and forget less useful signals over time. In this study, the LSTM model employs three layers, each with a hidden dimension of 64.
  • Temporal Convolutional Network (TCN) is a deep learning model that handles sequence data using 1D convolution layers. Unlike RNNs, it processes sequences in parallel and captures long-range dependencies using dilated convolutions. In this study, the TCN model consists of a single temporal block with 128 channels. Causal convolutions with a kernel size of 2 are used, and a dropout rate of 0.1 is applied.
  • Support Vector Regression (SVR) is a machine learning method that finds a line or curve to fit the data while keeping errors within a set margin. It works well on small to medium-sized datasets and handles both linear and non-linear trends. In this study, the SVR uses a linear kernel with the regularization parameter C set to 10.
  • Random Forest is an ensemble model that builds many decision trees and combines their results. It reduces overfitting and improves prediction accuracy by averaging the outcomes of individual trees. In this study, the Random Forest consists of 300 trees with a maximum tree depth of 30. The minimum number of samples required for node splitting and leaf nodes is set to 10 and 2, respectively.
  • XGBoost (Extreme Gradient Boosting) is a tree-based machine learning method built on the idea of gradient boosting. It constructs an ensemble of decision trees in a sequential way, where each new tree focuses on correcting the errors made by previous trees. The model uses gradient information from the loss function to guide the learning process, which helps improve prediction accuracy. In this study, the XGBoost model is configured with a squared error objective, 100 boosting trees, a learning rate of 0.1, and a maximum tree depth of 3.
  • Informer [55] is a Transformer-based model designed for long time-series forecasting. It introduces a probabilistic sparse self-attention mechanism that selects the most relevant queries, which reduces computational cost while keeping important temporal information. By combining sparse attention with an efficient encoder–decoder structure, Informer can handle long input sequences and is suitable for large-scale forecasting tasks. In this study, the Informer model adopts a hidden dimension of 512 with eight attention heads and consists of three encoder layers and two decoder layers. ProbSparse attention is used to improve computational efficiency, along with fixed time embeddings. The remaining settings follow standard configurations, including GELU activation and zero dropout.
@@ 3.5 Evaluation
Furthermore, to highlight the novelty of incorporating sentiments from both target companies and their related firms, we evaluate the impact of integrating sentiment information into a Transformer model for stock price forecasting, which will be illustrated in Section 4.3. For both the baseline scenario without sentiment and the experimental scenario with sentiment, we conduct ten independent runs per stock. This number is chosen as a common practice in empirical machine learning research to balance statistical robustness with computational efficiency. Furthermore, to examine the feature randomness and training variance, two experimental setups are considered: (i) fixing a single alpha set and repeating the experiment of all models for 10 runs to account for training variance; and (ii) fixing the training seed of the Transformer model and generating ten different alpha sets to account for feature randomness. The experimental setup (ii) includes the comparison between Transformer trained with and without sentiment information.
We report the average MSE and the standard deviation in each case. This ensures a fair and statistically sound comparison, allowing us to assess not only the average performance improvement of sentiment features, but also the consistency of the model predictions across multiple runs.
3.5.1 Diebold–Mariano Test
In addition to error-based metrics, we employ the Diebold–Mariano (DM) test [15] to assess whether the forecasting performance differences between models are statistically significant. Consider two competing forecasting models, denoted as model $A$ and model $B$. Let $y_{t}$ be the observed stock price at time $t$, and $\hat{y}_{t}^{A}$ and $\hat{y}_{t}^{B}$ be the corresponding out-of-sample forecasts. For a given loss function $L(\cdot)$, the loss differential is defined as
$d_{t}=L(y_{t},\hat{y}_{t}^{A})-L(y_{t},\hat{y}_{t}^{B}).$ (2)
In this study, we adopt the mean squared error as the loss function, i.e., $L(y_{t},\hat{y}_{t})=(y_{t}-\hat{y}_{t})^{2}$.
The DM test examines the null hypothesis that the two models have equal predictive accuracy, $H_{0}:\mathbb{E}[d_{t}]=0$. The test statistic is calculated as $S=\bar{d}/\sqrt{\hat{V}(\bar{d})}$, where $\bar{d}$ is the sample mean of the loss differential and $\hat{V}(\bar{d})$ is an estimate of its variance accounting for serial correlation.
To account for potential bias in small sample sizes and for the prediction horizon $h$, we apply the Harvey-Leybourne-Newbold (HLN) correction [18]. The corrected DM statistic, $S^{*}$, is computed as follows:
$S^{*}=S\cdot\sqrt{\frac{T+1-2h+h(h-1)/T}{T}},$ (3)
where $T$ represents the number of observations in the out-of-sample period and $h$ is the forecast horizon. In our empirical evaluation, we set $h=1$ for one-step-ahead predictions. The corrected statistic $S^{*}$ is then compared against the critical values of a Student’s $t$-distribution with (T1)(T-1) degrees of freedom. A statistically significant $S^{*}$ indicates that the difference in forecasting accuracy is unlikely to be driven by random variation.
Finally, since we perform multiple pairwise comparisons between different models, there is a higher risk of finding significant results by pure chance. To fix this, we apply the Holm-Bonferroni adjustment [19] to our $p$-values. This method works by ranking the $p$-values from smallest to largest and adjusting each one with a different weight. This ensures that our statistical conclusions remain strict and reliable even when many tests are conducted at the same time.
@@ 4 Results
Table 5 shows the LLM-generated formulaic alphas for all the five companies. Table 9 demonstrates the Transformer performance with and without sentiment scores as input. Figure 5 shows the reasoning process of the LLM to generate the formulaic alphas for Toyota. Table 6 shows the prediction performance in terms of MSE for all companies using the formulaic alphas in Table 5. Figure 6 visualizes the prediction performances of some models for HSBC stock prediction. Based on the figures, the models are able to capture stock price movements when LLM-generated features are used as inputs. More details of the results will be illustrated in the subsequent subsections.
@@ 4.2 Stock Prediction Performance
Table 6 reports the stock price prediction performance of different models measured by MSE. The comparison is conducted before inverse transformation, as our primary interest lies in the relative performance across models rather than the absolute price scale.
According to Table 6, for most of the time, Transformer has the best performance in terms of MSE, followed by Informer. The plotted graph for the Transformer (Figure 6) shows that the predicted stock prices closely follow the actual stock price trend, indicating its strong ability to capture underlying patterns in the data. In addition, the strong performance of the Transformer and Informer models highlights the advantages of attention mechanisms in sequential modeling.
@@
For traditional machine learning models, the SVR has the worst performance. The ensemble methods, random forest and XGBoost, can sometimes generate competitive performance compared to the Transformer (e.g., HSBC and Toyota).
Table 7 reports the overall Diebold–Mariano (DM) test results comparing the Transformer with the baseline models across all stocks. The statistical outcomes are consistent with the MSE rankings. In particular, the Transformer shows significantly better forecasting accuracy than LSTM, TCN, SVR, and the ensemble models, with all differences being statistically significant ($p<0.01$). In contrast, the difference between the Transformer and Informer is not statistically significant ($p>0.05$), suggesting that their similar attention-based structures lead to comparable predictive performance.
To provide a clearer view of the pairwise model comparisons, Table 8 reports the Diebold–Mariano test results for Pepsi stock. This stock is used as a representative example to illustrate the relative performance differences between models. The results show that the Transformer achieves statistically better forecasting accuracy than most baseline models. Similar patterns are observed for the other stocks in our experiments.
@@
[…]
@@
Table 6: MSE Comparison of Models for Stock Price Prediction with LLM-Generated Alphas, with lowest MSE in each column in bold and second-lowest underlined.
@@
XGBoost 0.0019 0.0005 0.0014 0.0011 0.0005
Table 7: Diebold–Mariano Test Summary: Transformer vs. Baseline Models Across All Stocks. Each cell reports the DM statistic with $p$-value in parentheses. A negative statistic indicates the Transformer’s superior predictive accuracy.
Baseline Model Apple HSBC Pepsi Tencent Toyota Informer 0.27-0.27 (0.787) 0.48-0.48 (0.631) 0.64-0.64 (0.524) 0.39-0.39 (0.697) 0.71-0.71 (0.478) LSTM 20.08-20.08*** 16.75-16.75*** 7.97-7.97*** 12.31-12.31*** 8.05-8.05*** TCN 29.63-29.63*** 12.56-12.56*** 10.62-10.62*** 16.26-16.26*** 9.88-9.88*** SVR 16.56-16.56*** 19.63-19.63*** 16.02-16.02*** 25.84-25.84*** 14.76-14.76*** Random Forest 8.44-8.44*** 13.27-13.27*** 5.34-5.34*** 10.53-10.53*** 5.12-5.12*** XGBoost 8.81-8.81*** 13.44-13.44*** 5.76-5.76*** 10.62-10.62*** 5.44-5.44***
Note: (***) denotes $p<0.01$. Negative DM values indicate lower loss for the Transformer model.
@@
Table 8: Diebold–Mariano test results for Pepsi stock. Each cell compares the row model against the column model. The test statistic is reported with the corresponding Holm–Bonferroni adjusted $p$-value shown in parentheses. The loss differential is defined as $d_{t}=L_{\text{row}}-L_{\text{column}}$, such that a negative and statistically significant DM statistic indicates superior predictive accuracy of the row model. Statistical significance is denoted by *** ($p<0.01$), ** ($p<0.05$), and * ($p<0.10$).
[…]